DIFFERENTIAL GEOMETRY/PDE SEMINAR

Friday, May 17, 2002 Padelford C-36 2:30 pm

The FBI transform and operators with rough symbols

Daniel Tataru (UC Berkeley)

The FBI transform, which has found applications in studying partial differential equations with limited regularity coefficients, is a continuous transform on the space of functions which simultaneously localises a function in the space variables as well as the frequency variables. It is roughly analagous to a continuous wavelet transform, but is better suited to hyperbolic equations, as well as problems in unique continuation. In this talk we review the basics of the FBI transform, as well as discuss recent applications for operators with nonsmooth symbols.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.